OpenStudy (anonymous):

MULTIPLE CHOICE An initial population of 175 quail increases at an annual rate of 22%. Write an exponential function to model the quail population. What will the approximate population be after 5 years? A. ƒ(x) = 175(0.22)^x; 473 B. ƒ(x) = (175 • 0.22)^x; 84,587,005 C. ƒ(x) = 175(22)^x; 901,885,600 D. ƒ(x) = 175(1.22)^x; 473

OpenStudy (anonymous):

using differential equations you should get something such as \[P=ce^{kt}\] at time 0, population = 175 P(0)=175 put this into the equation \[175=ce^{k(0)}\] \[175=c(1)=c\] so c = 175 now we have \[P=175e^{kt}\]

OpenStudy (anonymous):

oh wait this is not compounded it's a flat rate umm hold on

OpenStudy (anonymous):

so alright if you have an initial value 175, and you want to add 22 % each year. that'd mean after the first year you'd have 1.22. The only with 1.22 is your last answer

OpenStudy (anonymous):

Exponential equations have this general form: Y=a*(b^x) Making Y = Y(t), a = Y0, b = (1+k) and x = t, we have: Y(t)=Y0(1+k)^t, where k is the growth rate (in decimals) dand t is the period. Now try replacing the data and see where you get ;-)

OpenStudy (anonymous):

1.22 isn't an answer choice though?

OpenStudy (anonymous):

1+k=?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If k is the growth rate, k = 0.22, then...

OpenStudy (anonymous):

Now try to figure what is the value of Y(0) and t

OpenStudy (anonymous):

ok thanks